OpenStudy (anonymous):

Mike is six less than three times his daughter's age. Mike is 36; how old is his daughter?

OpenStudy (anonymous):

Working backwards, 36 + 6 = 42 divided by 3 = His daughters age.

OpenStudy (zepp):

Daughter's age * 3 -6 = Mike Mike is 36 Daughter's age * 3 - 6 = 36 Daughter's age * 3 = 36 + 6 Daughter's age = 42/3